OpenStudy (anonymous):

Is this right ? (2x-8)(2x+1)=x^2-16 2(x-4)(2x+1)=(x-4)(x+4) l /x-4 2(2x+1)=x+4 l -x+4 4x+2-x+4=0 3x+6=0 l -6 3x=-6 l /3 x=-2

OpenStudy (anonymous):

Yes .-.

OpenStudy (anonymous):

seems unlikely but maybe

hero (hero):

Following me around again @satellite73

OpenStudy (anonymous):

I got the same answer , lemme see if i can get a different answer by doing it right ..

OpenStudy (anonymous):

no it is not right

OpenStudy (anonymous):

@Hero yeah, i need to keep you honest!

OpenStudy (anonymous):

Here's how to figure out if you ever doubt yourself , take your number you get for x which in your case you got -2 , & simply substitute it for x .. so it should be (2)(-2)-8(2)(-2)+1 = -2 squared - 16

hero (hero):

I was hoping I would be able to travel to and fro without having to watch my back.

OpenStudy (anonymous):

What is it then ?

OpenStudy (debbieg):

The first problem is here: 2(x-4)(2x+1)=(x-4)(x+4) l /x-4 You can divide by x-4, if you don't know that x-4 is not=0.

hero (hero):

Use quad formula, complete the square or whatever method you feel comfortable with to isolate x.

OpenStudy (anonymous):

oops

OpenStudy (debbieg):

Compounded here: 2(2x+1)=x+4 l -x+4 4x+2-x+4=0 IF the step before were correct, you would subtract (x+4), not subtract x and add 4.

OpenStudy (anonymous):

\(4x^2-x^2=?\)

OpenStudy (debbieg):

*You CAN'T divide by x-4, if you don't know that x-4 is not=0.

OpenStudy (anonymous):

@Hero did i make you nervous or something?

hero (hero):

No, you didn't, I just goofed.

hero (hero):

(2x−8)(2x+1)=x^2−16 Multiply the right side: 2x(2x+1)−8(2x+1)=x^2−16 4x^2+2x−16x−8=x^2−16 4x^2−14x−8=x^2−16 Subtract x^2 from both sides: 3x^2−14x−8=−16 Add 16 to both sides: 3x^2−14x+8=0

hero (hero):

Now, you can use quad formula, or complete the square, or whatever method feels comfortable for you.

OpenStudy (anonymous):

I need to know what x is.

hero (hero):

Do you know any methods to isolate x such as factoring, quadratic formula or complete the square?

OpenStudy (anonymous):

I know methods to facrorise.

hero (hero):

Well, believe it or not 3x^2 -14x + 8 = 0 Is factorable

OpenStudy (anonymous):

a^2-2ab+b^2=(a-b)^2 ?

hero (hero):

That only works if you have a perfect square.

OpenStudy (anonymous):

Yes.

hero (hero):

3x^2 - 14x + 8 = 0 that's not a perfect square, but it is factorable. Try finding two numbers that multiply to get 24 yet add to get -14 m x n = 24 m + n = -14

OpenStudy (anonymous):

12 and 2.

hero (hero):

Actually -12 and -2

OpenStudy (anonymous):

Yes.

hero (hero):

We can replace -14 with -(12 + 2) in the equation: 3x^2 - (12 + 2)x + 8 = 0 3x^2 -12x - 2x + 8 = 0

hero (hero):

Now we can factor by grouping.

hero (hero):

If you factor the first two terms you get 3x(x - 4) If you factor the last two terms you gete -2(x - 4) So you have 3x(x - 4) - 2(x - 4) = 0 x - 4 is also common to both so you factor that out to get (x - 4)(3x - 2) = 0 From here, you use zero product property: x - 4 = 0 3x - 2 = 0 So there are two values of x